# Melody Is All You Need For Music Generation

### v3

@@ Abstract.
We present the Melody Guided Music Generation ($\textbf{MG}^{2}$) model, a novel approach using melody to guide the text-to-music generation that, despite a pretty simple method and extremely limited resources, achieves excellent performance. Specifically, we first align the text with audio waveforms and their associated melodies using the newly proposed Contrastive Language-Music Pretraining, enabling the learned text representation fused with implicit melody information. Subsequently, we condition the retrieval-augmented diffusion module on both text prompt and retrieved melody. This allows $\text{MG}^{2}$ to generate music that reflects the content of the given text description, meantime keeping the intrinsic harmony under the guidance of explicit melody information. We conducted extensive experiments on two public datasets: MusicCaps and MusicBench. The experimental results demonstrate that the proposed $\text{MG}^{2}$ model surpasses current open-source text-to-music generation models, utilizing fewer than 1/3 of the parameters and less than 1/200 of the training data compared to state-of-the-art counterparts. Furthermore, we carried out comprehensive human evaluations to explore the potential applications of $\text{MG}^{2}$ in real-world scenarios [^1] We release the code and dataset for reproducing in GitHub: https://github.com/shaopengw/Awesome-Music-Generation. What’s more, $\text{MG}^{2}$ is being integrated in the Hugging Face Transformers library (See more here)..
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Applied computing Sound and music computing
@@ 1. Introduction
Guided by the melody, the generated music not only captures the semantic content of input text but also mitigates monotony, noise, and disjointed melody, ensuring harmony across and within segments. Extensive experiments demonstrate that the proposed $\text{MG}^{2}$, trained on MusicCaps and MusicBench—two public datasets totaling approximately 132 hours of music and utilizing 416 million parameters, outperforms state-of-the-art open-source models trained with over 200 times the data or 3 times the parameters.
Moreover, music generation emphasizes the artistry of creation, distinguishing it from typical text or image generation. This artistic quality cannot be adequately assessed using objective metrics alone, necessitating subjective evaluation. However, most previous studies (Liu et al., 2023, 2024a; Copet et al., 2024) have not thoroughly or accurately assessed generated music, often relying on fewer than 10 participants and simplistic evaluation criteria. In contrast, we assess the generated music across dimensions of recognizability, text relevance, satisfaction, quality, and market potential by engaging 135 general users, 18 musicians, and 20 short-video creators, using newly designed, comprehensive questionnaires. The results underscore the potential commercial value of $\text{MG}^{2}$ in real-world applications.
@@ 2.1. Text-to-Music Generation
Text-to-Music (TTM) aims to generate music based on a given text description, forming a specialized sub-domain of Text-to-Audio (TTA) generation (Liu et al., 2023; Kreuk et al., 2022; Yang et al., 2023). Unlike TTA, TTM requires both semantic alignment with the input description and harmonious melody in the generated music. Current research primarily follows two technical approaches: generating musical scores (Agostinelli et al., 2023; Yuan et al., 2024a; Ding et al., 2024) and generating audio files (Copet et al., 2024; Chen et al., 2024b; Lam et al., 2024). The first approach treats music as a structured language, generating it under a sequence-to-sequence framework. For example, ChatMusician (Yuan et al., 2024a) employs ABC notation to produce and interpret symbolic music, while SongComposer (Ding et al., 2024) generates lyrics and melodies in a tuple format, aligning lyrics with musical note attributes. The second approach aims to produce playable music by directly generating audio files that capture the original sound information. For example, AudioLDM (Liu et al., 2023) and TANGO (Ghosal et al., 2023) use text prompts as guidance to control the audio generation process. To better capture the attributes of music, some works incorporate information on rhythm or melody from various perspectives. For instance, Music ControlNet (Wu et al., 2024) proposed a method that leverages both input text and time-varying attributes—including melody, dynamics, and rhythm—as control signals to guide music generation. Similarly, Mustango (Melechovsky et al., 2023) utilizes text alongside generated beats and chords to guide the music generation process. Specifically, the beats and chords are derived from the input text using a trained DeBERTa Large model.
@@ 2.2. Generative Models
In recent years, several types of deep neural network-based music generation methods have been developed, including RNN-based models (Boulanger-Lewandowski et al., 2012; Waite et al., 2016), VAE-based models (Roberts et al., 2018; Wang et al., 2020), GAN-based models (Yang et al., 2017; Donahue et al., 2019b) and transformer-based models (Donahue et al., 2019a; Huang and Yang, 2020). For example, GANSynth (Engel et al., 2019) utilized GANs to generate high-fidelity, locally coherent audio. More recently, diffusion models have become the mainstream approach for music generation. Diffusion models (Nichol and Dhariwal, 2021; Song et al., 2020) operate by learning a generative process through a forward step that progressively adds noise to a sample and a reverse step that aims to denoise it. These models have gained significant attention and have been successfully applied to various generation tasks, including image generation (Zhang et al., 2023; Saharia et al., 2022; Dhariwal and Nichol, 2021), text generation (Chen et al., 2024a; Wu et al., 2023b; Gong et al., 2022), video generation (Liu et al., 2024b), and music generation (Chen et al., 2024b; Huang et al., 2023; Bai et al., 2024). For instance, AudioLDM (Liu et al., 2023) employs CLAP (Wu et al., 2023a) to align text and audio, using the aligned audio as a condition for the diffusion model to generate latent audio representations during training and text as the condition during inference. In contrast, Seed-Music (Bai et al., 2024) integrates both auto-regressive language modeling with diffusion approaches, enabling the model to handle both music generation and music editing tasks.
@@ 2.3. Retrieval-augmented Generation
Retrieval-augmented Generation (RAG) enhances the generation process by integrating external knowledge related to the generated content. This approach improves large language models in multiple aspects, such as reducing hallucination (Shuster et al., 2021), safeguarding privacy (Zeng et al., 2024), and enhancing the ability to answer questions with long context (Edge et al., 2024), making it widely adopted. RAG has also been applied in image generation. For instance, Blattmann et al. (Blattmann et al., 2022) proposed constructing an independent image database to retrieve a set of nearest neighbors as conditions for the generative model. Similarly, KNN-diffusion (Sheynin et al., 2022) introduced a technique to train a relatively small text-to-image diffusion model without text data, enabling out-of-distribution image generation by swapping the retrieval database during inference. In the field of audio generation, Yuan et al. (Yuan et al., 2024b) proposed using retrieved text-audio pairs, given a text prompt as the query, to perform text-to-audio generation.
@@ 5.1. Dataset
We evaluate all models using the well-known open-access music datasets, MusicCaps (Agostinelli et al., 2023) and MusicBench (Melechovsky et al., 2023).
@@ 5.2. Baselines
  • Mustango (Melechovsky et al., 2023): This work introduces a Music-Domain-Knowledge-Informed UNet guidance module to direct the generation of music, effectively capturing music-specific factors for text-to-music generation.
@@ 5.3. Experimental Setting
We implement all models using the official codes for the baseline models. For the CLMP of $\text{MG}^{2}$, we choose Adam (Kingma, 2014) as the optimizer, setting the learning rate to $1\times 10^{-5}$, the batch size to 48, and the number of epochs to 90. For the retrieval-augmented diffusion module and the decoding module of $\text{MG}^{2}$, we adopt AdamW (Loshchilov and Hutter, 2017) as the optimizer, with a learning rate of $1\times 10^{-4}$, a batch size of 96, a denoising step of 100, an unconditional guide weight of 3, a compression level of 4, and a total of 60,000 training steps. The CLMP is trained on one NVIDIA RTX 4090 24GB GPU, while the retrieval-augmented diffusion module is trained on one NVIDIA A800 80GB GPU. We use the officially released checkpoint of AudioLDM-M and fine-tune it on the MusicCaps and MusicBench datasets, respectively. During training, we first use audio as the query to develop the model’s ability to generate music, then use text as the query to enhance its ability to understand the semantic information of text input. For TANGO, we refer to the results tested by Mustango. For Mustango, since it is fine-tuned on MusicCaps and MusicBench, we directly test it using the official checkpoint on the same test set as $\text{MG}^{2}$. The training dataset of AudioLDM 2 includes AudioSet, which incorporates MusicCaps and MusicBench. Therefore, we also directly test it without fine-tuning.
@@ 5.4. Evaluation Protocol
5.5. Music Generation Results
As shown in Table 1, the proposed $\text{MG}^{2}$ outperforms all baselines in terms of FAD and KL divergence. Specifically, $\text{MG}^{2}$ exceeds the performance of the strongest baseline, AudioLDM-Full, by 1.30 on MusicCaps and by 0.75 on MusicBench based on the FAD indicator. Notably, $\text{MG}^{2}$ is trained on MusicCaps with 5,000 music tracks and MusicBench with 50,000 tracks, totaling approximately 132 hours of audio. In contrast, AudioLDM-Full is trained on 29,510 hours of samples—about 200 times more than the data used by $\text{MG}^{2}$. Furthermore, $\text{MG}^{2}$ achieves these results with only 1/3 of the parameters compared to the strong baseline, Mustango. Additionally, due to the overlap in fine-tuning datasets (i.e., MusicCaps and AudioCaps), models like TANGO and AudioLDM may have been exposed to portions of the MusicCaps test set in our experimental setup. Despite this, $\text{MG}^{2}$ still achieves comparable performance on the IS indicator, further validating its effectiveness. We attribute these improvements to the implicit and explicit use of melody guidance, which ensures that the music generated by $\text{MG}^{2}$ is both harmonious across fragments and within each individual segment, resulting in less noise and more fitting rhythm.
@@ 5.6. Multimodal Alignment
Additionally, we demonstrate the effectiveness of alignment by using both audio and text from the same pair to guide the music generation process. As illustrated in Figure 2, the music generated with text guidance and audio guidance results in nearly identical music spectrograms in each sub-figure. It is important to note that both the audio and text representations are derived from the CLMP. This alignment is due to the semantic correspondence between the text modality and the waveform modality. On one hand, the results indicate that the text representation and audio waveform representation from the same pair are highly similar. On the other hand, the retrieved melody guidance is also consistent, reflecting the high similarity between the two input queries. In conclusion, the proposed three-modality alignment achieves excellent performance not only in experimental metrics but also in the real music generation process.
@@ 6. Human Evaluation
We designed questionnaires with demographic questions and specific evaluation criteria for three user groups to comprehensively assess the effectiveness of the proposed $\text{MG}^{2}$ model. These criteria cover four key aspects: recognizability, text relevance, satisfaction, quality, and market potential (see Appendix A for details). Each user group provides unique insights based on their expertise: General users assess music based on subjective satisfaction, reflecting personal enjoyment but lacking the technical perspective for professional quality evaluation. Musicians offer professional insights into music quality but may not gauge the business potential. Short video creators evaluate market potential, leveraging their experience in content creation. We included seven sample pieces (Table 4) and made them accessible on our Website.
In the demographic section, we also gathered each of the 163 subjects’ attitudes toward AI-generated music, displayed in Figure 7. Results indicate that 88.4% of participants are supportive or neutral about AI music, suggesting a promising future for AI in music generation.
@@ 6.1. Recognizability
We then asked all 163 subjects to identify AI-generated music from a set of five music clips, which included two pieces created by $\text{MG}^{2}$ and three human-made compositions. Participants could select none, one, or multiple options. Table 3 on our Website provides these music examples for reference. As illustrated in Figure 8, only 6.7% of participants successfully identified both AI-generated pieces among the five clips. Over half mistakenly labeled all human-made music as AI-generated, failing to identify any of the AI-generated pieces. This result contrasts with AI creations in image and text domains, where people often identify AI content despite high model quality and few visible flaws. Further analysis of the selections shows that the AI-generated pieces (Music 2 and Music 4) received the fewest ’AI music’ votes, while human-made tracks, such as Music 5 and Music 1, attracted significantly more. This suggests that $\text{MG}^{2}$ effectively produces music that many listeners perceive as human-created, underscoring its high quality and realistic output.

### v4

# Melody-Guided Music Generation

@@ Abstract.
We present the Melody-Guided Music Generation ($\textbf{MG}^{2}$) model, a novel approach using melody to guide the text-to-music generation that, despite a simple method and limited resources, achieves excellent performance. Specifically, we first align the text with audio waveforms and their associated melodies using the newly proposed Contrastive Language-Music Pretraining, enabling the learned text representation fused with implicit melody information. Subsequently, we condition the retrieval-augmented diffusion module on both text prompt and retrieved melody. This allows $\text{MG}^{2}$ to generate music that reflects the content of the given text description, meantime keeping the intrinsic harmony under the guidance of explicit melody information. We conducted extensive experiments on two public datasets: MusicCaps and MusicBench. Surprisingly, the experimental results demonstrate that the proposed $\text{MG}^{2}$ model surpasses current open-source text-to-music generation models, achieving this with fewer than 1/3 of the parameters or less than 1/200 of the training data compared to state-of-the-art counterparts. Furthermore, we conducted comprehensive human evaluations involving three types of users and five perspectives, using newly designed questionnaires to explore the potential real-world applications of $\text{MG}^{2}$  [^1] We release the code and dataset for reproducing in GitHub: https://github.com/shaopengw/Awesome-Music-Generation. What’s more, $\text{MG}^{2}$ is being integrated in the Hugging Face Transformers library (See more here)..
@@
ccs: Applied computing Sound and music computing
@@ 1. Introduction
Guided by the melody, the generated music not only captures the semantic content of input text but also mitigates monotony, noise, and disjointed melody, ensuring harmony across and within segments. Extensive experiments demonstrate that the proposed $\text{MG}^{2}$, trained on MusicCaps and MusicBench—two public datasets totaling approximately 132 hours of music and utilizing 416 million parameters, outperforms state-of-the-art open-source models trained with over 200 times the data or 8 times the parameters.
Moreover, music generation emphasizes the artistry of creation, distinguishing it from typical text or image generation. This artistic quality cannot be adequately assessed using objective metrics alone, necessitating subjective evaluation. However, most previous studies (Liu et al., 2023, 2024a; Copet et al., 2023) have not thoroughly or accurately assessed generated music, often relying on fewer than 10 participants and simplistic evaluation criteria. In contrast, we assess the generated music across dimensions of recognizability, text relevance, satisfaction, quality, and market potential by engaging 124 general users, 18 musicians, and 20 short-video creators, using newly designed, comprehensive questionnaires. The results underscore the potential commercial value of $\text{MG}^{2}$ in real-world applications.
@@ 2.1. Text-to-Music Generation
Text-to-Music (TTM) aims to generate music based on a given text description, forming a specialized sub-domain of Text-to-Audio (TTA) generation (Liu et al., 2023; Kreuk et al., 2023; Yang et al., 2023). Unlike TTA, TTM requires both semantic alignment with the input description and harmonious melody in the generated music. Current research primarily follows two technical approaches: generating musical scores (Agostinelli et al., 2023; Yuan et al., 2024a; Ding et al., 2024) and generating audio files (Copet et al., 2023; Chen et al., 2024b; Lam et al., 2024). The first approach treats music as a structured language, generating it under a sequence-to-sequence framework. For example, ChatMusician (Yuan et al., 2024a) employs ABC notation to produce and interpret symbolic music, while SongComposer (Ding et al., 2024) generates lyrics and melodies in a tuple format, aligning lyrics with musical note attributes. The second approach aims to produce playable music by directly generating audio files that capture the original sound information. For example, AudioLDM (Liu et al., 2023) and TANGO (Ghosal et al., 2023) use text prompts as guidance to control the audio generation process. To better capture the attributes of music, some works incorporate information on rhythm or melody from various perspectives. For instance, Music ControlNet (Wu et al., 2024) proposed a method that leverages both input text and time-varying attributes—including melody, dynamics, and rhythm—as control signals to guide music generation. Similarly, Mustango (Melechovsky et al., 2024) utilizes text alongside generated beats and chords to guide the music generation process. Specifically, the beats and chords are derived from the input text using a trained DeBERTa Large model.
@@ 2.2. Generative Models
In recent years, several types of deep neural network-based music generation methods have been developed, including RNN-based models (Boulanger-Lewandowski et al., 2012; Waite et al., 2016), VAE-based models (Roberts et al., 2018; Wang et al., 2020), GAN-based models (Yang et al., 2017; Donahue et al., 2019) and transformer-based models (Donahue et al., [n. d.]; Huang and Yang, 2020). For example, GANSynth (Engel et al., 2019) utilized GANs to generate high-fidelity, locally coherent audio. More recently, diffusion models have become the mainstream approach for music generation. Diffusion models (Nichol and Dhariwal, 2021; Song et al., 2020) operate by learning a generative process through a forward step that progressively adds noise to a sample and a reverse step that aims to denoise it. These models have gained significant attention and have been successfully applied to various generation tasks, including image generation (Zhang et al., 2023; Saharia et al., 2022; Dhariwal and Nichol, 2021), text generation (Chen et al., 2024a; Wu et al., 2023b; Gong et al., 2023), video generation (Liu et al., 2024b), and music generation (Chen et al., 2024b; Huang et al., 2023; Bai et al., 2024). For instance, AudioLDM (Liu et al., 2023) employs CLAP (Wu et al., 2023a) to align text and audio, using the aligned audio as a condition for the diffusion model to generate latent audio representations during training and text as the condition during inference. In contrast, Seed-Music (Bai et al., 2024) integrates both auto-regressive language modeling with diffusion approaches, enabling the model to handle both music generation and music editing tasks.
@@ 2.3. Retrieval-augmented Generation
Retrieval-augmented Generation (RAG) enhances the generation process by integrating external knowledge related to the generated content. This approach improves large language models in multiple aspects, such as reducing hallucination (Shuster et al., 2021), safeguarding privacy (Zeng et al., 2024), and enhancing the ability to answer questions with long context (Edge et al., 2024), making it widely adopted. RAG has also been applied in image generation. For instance, Blattmann et al. (Blattmann et al., 2022) proposed constructing an independent image database to retrieve a set of nearest neighbors as conditions for the generative model. Similarly, KNN-diffusion (Sheynin et al., 2023) introduced a technique to train a relatively small text-to-image diffusion model without text data, enabling out-of-distribution image generation by swapping the retrieval database during inference. In the field of audio generation, Yuan et al. (Yuan et al., 2024b) proposed using retrieved text-audio pairs, given a text prompt as the query, to perform text-to-audio generation.
@@ 5.1. Dataset
We evaluate all models using the well-known open-access music datasets, MusicCaps (Agostinelli et al., 2023) and MusicBench (Melechovsky et al., 2024).
@@ 5.2. Baselines
  • MusicLM (Agostinelli et al., 2023): This work models the conditional music generation as a hierarchical sequence-to-sequence task.
@@
  • MusicGen (Copet et al., 2023): This work proposes to use a single-stage transformer LM together with codebook interleaving patterns for conditional music generation.
@@
  • Mustango (Melechovsky et al., 2024): This work introduces a Music-Domain-Knowledge-Informed UNet guidance module to direct the generation of music, effectively capturing music-specific factors for text-to-music generation.
  • FluxMusic (Fei et al., 2024): This work proposes a diffusion-based rectified flow transformers for text-to-music generation.
@@ 5.3. Experimental Setting
We implement all models using the official codes for the baseline models. For the CLMP of $\text{MG}^{2}$, we choose Adam (Kingma, 2014) as the optimizer, setting the learning rate to $1\times 10^{-5}$, the batch size to 48, and the number of epochs to 90. For the retrieval-augmented diffusion module and the decoding module of $\text{MG}^{2}$, we adopt AdamW (Loshchilov and Hutter, 2017) as the optimizer, with a learning rate of $1\times 10^{-4}$, a batch size of 96, a denoising step of 100, an unconditional guide weight of 3, a compression level of 4, and a total of 60,000 training steps. During training, we first use audio as the query to develop the model’s ability to generate music, then use text as the query to enhance its ability to understand the semantic information of text input. The CLMP is trained on one NVIDIA RTX 4090 24GB GPU, while the retrieval-augmented diffusion module is trained on one NVIDIA A800 80GB GPU. Since there is no official code or model checkpoint for MusicLM, we use the reported results from the MusicLM paper. We use the officially released checkpoint of AudioLDM-M and fine-tune it on the MusicCaps and MusicBench datasets, respectively. For TANGO, we refer to the results tested by Mustango. The training dataset of AudioLDM 2 includes AudioSet, which incorporates MusicCaps and MusicBench. Therefore, we directly test it without fine-tuning. For Mustango, since it is fine-tuned on MusicCaps and MusicBench, we also directly test it using the official checkpoint on the same test set as $\text{MG}^{2}$. For MusicGen and FluxMusic, we use the official code and corresponding largest checkpoints to perform inferences without fine-tuning, following the procedures described in their respective papers.
@@ 5.4. Evaluation Protocol
5.5. Music Generation
Main Results. As shown in Table 1, the proposed $\text{MG}^{2}$ outperforms all baselines in terms of FAD and KL divergence. Specifically, $\text{MG}^{2}$ exceeds the performance of the strongest baseline, AudioLDM-Full, by 1.30 on MusicCaps and by 0.75 on MusicBench based on the FAD indicator. Notably, $\text{MG}^{2}$ is trained on MusicCaps with 5,000 music tracks and MusicBench with 50,000 tracks, totaling approximately 132 hours of audio. In contrast, AudioLDM-Full is trained on 29,510 hours of samples—about 200 times more than the data used by $\text{MG}^{2}$. Furthermore, $\text{MG}^{2}$ achieves these results with only 1/3 of the parameters compared to the strong baseline, Mustango. Compared to other baselines such as MusicLM, MusicGen and FluxMusic, the gap in training resources and model performance grows larger. Specifically, $\text{MG}^{2}$ achieves better performance in terms of FAD and KL indicators using only 1/8 of the parameters of MusicGen, 1/5 of the parameters of FluxMusic, and less than 1/2000 of the training data used by MusicLM. Additionally, due to the overlap in fine-tuning datasets (i.e., MusicCaps and AudioCaps), models like TANGO and AudioLDM may have been exposed to portions of the MusicCaps test set in our experimental setup. Despite this, $\text{MG}^{2}$ still achieves comparable performance on the IS indicator, further validating its effectiveness.
Analysis. Basically, these models can be divided into two types based on the generation module: transformer-based models, including MusicLM and MusicGen, and diffusion-based models, which include the remaining models. Compared to diffusion-based models, transformer-based models require more training data to achieve comparable or better performance. This is likely because transformer-based models need large amounts of data to develop emergent abilities, similar to language models. While transformer-based models excel at modeling temporal information, they struggle with capturing fine-grained patterns, as music is inherently continuous. AudioLDM, TANGO, AudioLDM 2, and FluxMusic follow a similar approach by using text representations to condition the diffusion process. This ensures that the semantic information of the generated music is preserved, and this capability can be further enhanced with more training data, as demonstrated by comparisons between AudioLDM 2-Music and AudioLDM 2-Full, as well as FluxMusic. Nonetheless, these models still struggle to generate harmonious rhythm. As a result, melody has played a crucial role in music generation, as evidenced by MusicLM, MusicGen, and Mustango. While both MusicLM and MusicGen use melody solely as an input to guide the generation process, they lack explicit melody guidance for text inputs. Mustango addresses this by predicting melody information from the text input. However, this approach does not provide additional information theoretically, as the amount of input data remains unchanged. In contrast, $\text{MG}^{2}$ outperforms all these baselines overall, using fewer model parameters and less training data. Specifically, the melody vector database, serving as an external knowledge base, combined with the diffusion module, makes $\text{MG}^{2}$ highly efficient, relying on fewer parameters and less training data. The implicit and explicit use of melody guidance ensures that the music generated by $\text{MG}^{2}$ is harmonious both across fragments and within individual segments, resulting in reduced noise and a more fitting rhythm.
@@ 5.6. Multimodal Alignment
Additionally, we demonstrate the effectiveness of alignment by using both audio waveforms and text descriptions from the same pairs, randomly chosen from the MusicBench dataset, to guide the music generation process. As illustrated in Figure 2, the music generated with text input and waveform input results in nearly identical music spectrograms in each sub-figure. Due to space limitations, we provide the text descriptions of these pairs in Appendix A. It is important to note that both the audio and text representations are derived from the CLMP. This alignment is due to the semantic correspondence between the text modality and the waveform modality. On one hand, the results indicate that the text representation and audio waveform representation from the same pair are highly similar. On the other hand, the retrieved melody guidance is also consistent, reflecting the high similarity between the two input queries. In conclusion, the proposed three-modality alignment achieves excellent performance not only in experimental metrics but also in the real music generation process.
@@ 6. Human Evaluation
We designed questionnaires with demographic questions and specific evaluation criteria for three user groups to comprehensively assess the effectiveness of the proposed $\text{MG}^{2}$ model. These criteria cover four key aspects: recognizability, text relevance, satisfaction, quality, and market potential (see Appendix B for details). Each user group provides unique insights based on their expertise: General users assess music based on subjective satisfaction, reflecting personal enjoyment but lacking the technical perspective for professional quality evaluation. Musicians offer professional insights into music quality but may not gauge the business potential. Short video creators evaluate market potential, leveraging their experience in content creation. We included seven sample pieces (Table 4) and made them accessible on our Website.
In the demographic section, we also gathered each of the 162 subjects’ attitudes toward AI-generated music, displayed in Figure 7. Results indicate that 88.4% of participants are supportive or neutral about AI music, suggesting a promising future for AI in music generation.
@@ 6.1. Recognizability
We then asked all 162 subjects to identify AI-generated music from a set of five music clips, which included two pieces created by $\text{MG}^{2}$ and three human-made compositions. Participants could select none, one, or multiple options. Table 3 on our Website provides these music examples for reference. As illustrated in Figure 8, only 6.7% of participants successfully identified both AI-generated pieces among the five clips. Over half mistakenly labeled all human-made music as AI-generated, failing to identify any of the AI-generated pieces. This result contrasts with AI creations in image and text domains, where people often identify AI content despite high model quality and few visible flaws. Further analysis of the selections shows that the AI-generated pieces (Music 2 and Music 4) received the fewest “AI music” votes, while human-made tracks, such as Music 5 and Music 1, attracted significantly more. This suggests that $\text{MG}^{2}$ effectively produces music that many listeners perceive as human-created, underscoring its high quality and realistic output.